1 The Greek language

The file greek.dtx^{1} defines all the language definition macros for the Greek language.

The macro \LdfInit takes care of preventing that this file is loaded more than once, checking the category code of the @ sign, etc.

1 ⟨∗code⟩

 2 \LdfInit{greek}\captionsgreek

When this file is read as an option, i.e. by the \usepackage command, greek could be an ‘unknown’ language in which case we have to make it known. So we check for the existence of \l@greek to see whether we have to do something here.

 3 \ifx\l@greek\@undefined
 4 \@nopatterns{greek}
 5 \adddialect\l@greek0\fi

Tyesetting Greek texts implies that a special set of fonts needs to be used. The current support for greek uses a set of fonts that more or less ‘match’ with the computer modern fonts. Therefore we define the Local GReek encoding (LGR, see the file greek.fdd). We make sure that this encoding is known to L^{A}TEX.

 6 \input{LGRenc.def}

\latinencoding We need to know the encoding for text that is supposed to be typeset in latin text. We assume that it will be the encoding which is active at the end of the babel package. If the fontenc package is loaded later, then... too bad!

 7 \AtEndOfPackage{\edef\latinencoding{\cf@encoding}}

Now we define two commands that offer the possibility to switch between greek and roman encodings.

\greektext \latintext The command \greektext will switch from latin font encoding to the greek font encoding, the command \latintext switches back. This assumes that the ‘normal’ font encoding is a latin one. These commands are declarations, for shorter peaces of text the commands \textlatin and \textgreek can be used.

 8 \DeclareRobustCommand{\greektext}{%
 9 \fontencoding{LGR}\selectfont
10 \def\encodingdefault{LGR}}
11 \DeclareRobustCommand{\latintext}{%
12 \fontencoding{\latinencoding}\selectfont
13 \def\encodingdefault{\latinencoding}}

\textgreek \textlatin These commands take an argument which is then typeset using the requested font encoding. In order to avoid many encoding switches both commands operate in a local scope.

14 \DeclareRobustCommand{\textgreek}[1]{{\greektext #1}}
15 \DeclareRobustCommand{\textlatin}[1]{{\latintext #1}}

\textol A last aspect of the set of fonts provided with this version of support for typesetting Greek texts is that it contains an outline font. In order to make it available we define the command \textol.

1The file described in this section has version number v1.0b and was last revised on 1996/12/23. The original author is Apostolos Syropoulos (apostolo@platon.ee.duth.gr).

16 \def\outlfamily{\usefont{LGR}{cmro}{m}{n}}
17 \DeclareTextFontCommand{\textol}{\outlfamily}

The next step consists of defining commands to switch to (and from) the Greek language.

\greekhyphenmins This macro is used to store the correct values of the hyphenation parameters

\lefthyphenmin and \righthyphenmin.
18 \def\greekhyphenmins{23}

\captionsgreek The macro \captionsgreek defines all strings used in the four standard documentclasses provided with L^{A}TEX.

19 \addto\captionsgreek{%
20 \def\prefacename{Pr'ologoc}%
21 \def\refname{Anafor'ec}%
22 \def\abstractname{Per'ilhyh}%
23 \def\bibname{Bibliograf'ia}%
24 \def\chaptername{Kef'alaio}%
25 \def\appendixname{Par'arthma}%
26 \def\contentsname{Perieq'omena}%
27 \def\listfigurename{Kat'alogoc Sqhm'atwn}%
28 \def\listtablename{Kat'alogoc Pin'akwn}%
29 \def\indexname{Euret'hrio}%
30 \def\figurename{Sq'hma}%
31 \def\tablename{P'inakac}%
32 \def\partname{M'eroc}%
33 \def\enclname{Sunhmm'ena}%
34 \def\ccname{Koinopo'ihsh}%
35 \def\headtoname{Proc}%
36 \def\pagename{Sel'ida}%
37 \def\seename{bl'epe}%
38 \def\alsoname{bl'pe ep'ishc}%

39 }

\dategreek The macro \dategreek redefines the command \today to produce greek dates.

40 \def\dategreek{%
41 \def\today{\number\day \space%
42  \ifcase\month\or
43  Ianouar'iou\or Febrouar'iou\or Mart'iou\or April'iou\or
44  Ma'"iou\or Ioun'iou\or Ioul'iou\or Augo'ustou\or
45  Septembr'iou\or Oktobr'iou\or Noembr'iou\or Dekembr'iou\fi
46  \space \number\year}}

\extrasgreek \noextrasgreek The macro \extrasgreek will perform all the extra definitions needed for the Greek language. The macro \noextrasgreek is used to cancel the actions of \extrasgreek. For the moment these macros switch the fontencoding used and the definition of the internal macros \@alph and \@Alph because the greek alpabetisation differs considerably from the roman alpabetisation..

47 \addto\extrasgreek{\greektext}
48 \addto\noextrasgreek{\latintext}

\greek@alph \greek@Alph With the latin transcription used to create documents that will be typeset in the greek language we need to adopt the alphabetisation considerably. Therefore we have to redefine the internal L^{A}TEX commands \@alph and \@Alph.

We need to be able to switch between the original definitions and the greek ones, so we first ‘save’ the orginal definitions.

49 \let\latin@alph\@alph
50 \let\latin@Alph\@Alph

Then we define the greek versions

51 \def\greek@alph#1{%
52 \ifcase #1\or a\or b\or g\or d\or e\or st\or z\or h\or j\or i\or
53  ia\or ib\or ig\or id\or ie\or ist\or iz\or ih\or ij\or k\or ka\or
54  kb\or kg\or \kd \or ke\or kst
55 \else
56  \@ctrerr
57 \fi
58 $'$}
59 \def\greek@Alph#1{%
60 \ifcase #1\or A\or B\or G\or D\or E\or St\or Z\or H\or J\or I\or
61  IA\or IB\or IG\or ID\or IE\or IST\or IZ\or IH\or IJ\or K\or KA\or
62  KB\or KG\or \KD \or KE\or KST
63 \else
64  \@ctrerr
65 \fi
66 $'$}

Now we can set up the switching.

67 \addto\extrasgreek{%
68 \let\@alph\@greek@alph
69 \let\@Alph\@greek@Alph}
70 \addto\noextrasgreek{%
71 \let\@alph\latin@alph
72 \let\@Alph\latin@Alph}

\greek@roman \greek@Roman To prevent roman numerals being typeset in greek letters we need to adopt the internal L^{A}TEX commands \@roman and \@Roman.

73 \let\latin@roman\@roman
74 \let\latin@Roman\@Roman
75 \def\greek@roman#1{\textlatin{\latin@roman{#1}}}
76 \def\greek@Roman#1{\textlatin{\latin@Roman{#1}}}
77 \addto\extrasgreek{%
78 \let\@roman\@greek@roman
79 \let\@Roman\@greek@Roman}
80 \addto\noextrasgreek{%
81 \let\@roman\latin@roman
82 \let\@Roman\latin@Roman}

We provide access to a few extra greek characters. They are only available in one particular font, therefore we first define a ‘helper’ macro to select the correct font.

83 \def\greek@char#1{{%
84  \fontfamily\rmdefault
85  \fontseries\mddefault
86  \fontshape\scdefault
87  \selectfont\char#1}}
88 \DeclareTextCommand{\tao}{LGR}{\greek@char{"7F}}
89 \DeclareTextCommand{\Qoppa}{LGR}{\greek@char{"43}}

90 \DeclareTextCommand{\qoppa}{LGR}{\greek@char{"5B}}
91 \DeclareTextCommand{\varqoppa}{LGR}{\greek@char{"5C}}
92 \DeclareTextCommand{\Sampi}{LGR}{\greek@char{"5F}}
93 \DeclareTextCommand{\sampi}{LGR}{\greek@char{"5E}}
94 \DeclareTextCommand{\Digamma}{LGR}{\greek@char{"17}}

The amssymb package defines a \digamma command, so in order to avoid problems we spell here digamma intentionally erroneous with a double d.

95 \DeclareTextCommand{\ddigamma}{LGR}{\greek@char{"60}}
96 \DeclareTextCommand{\vardigamma}{LGR}{\greek@char{"5D}}

Now make sure that these commands can also be used outside of the greek font encoding.

97 \ProvideTextCommandDefault{\tao}{\UseTextSymbol{LGR}{\tao}}
98 \ProvideTextCommandDefault{\Qoppa}{\UseTextSymbol{LGR}{\Qoppa}}
99 \ProvideTextCommandDefault{\qoppa}{\UseTextSymbol{LGR}{\qoppa}}
100 \ProvideTextCommandDefault{\varqoppa}{\UseTextSymbol{LGR}{\varqoppa}}
101 \ProvideTextCommandDefault{\Sampi}{\UseTextSymbol{LGR}{\Sampi}}
102 \ProvideTextCommandDefault{\sampi}{\UseTextSymbol{LGR}{\sampi}}
103 \ProvideTextCommandDefault{\Digamma}{\UseTextSymbol{LGR}{\Digamma}}
104 \ProvideTextCommandDefault{\ddigamma}{\UseTextSymbol{LGR}{\ddigamma}}
105 \ProvideTextCommandDefault{\vardigamma}{%
106 \UseTextSymbol{LGR}{\vardigamma}}

The macro \ldf@finish takes care of looking for a configuration file, setting the main language to be switched on at \begin{document} and resetting the category code of @ to its original value.

107 \ldf@finish{greek}

108 ⟨/code⟩